STATISTICS Quiz 4 1. Southern California Edison Co. has begun to promote ene discounted rates to consumers who k certain established eep their energy usage below ubsidy standards. A reduced their electricity cent report notes that 70% of Catalina Island residents have usage sufficiently to qualify for discounted rates. If 5 residential re su subscribers are randomly selected from Catalina Island a) find the probability ht al iv quall or the favorable ratos: (3 pts.) b) find the probability that at least four qualify for the favorable rates. (3 pts.) c) find the probability that most of the residents qualify for the favorable rates. 24,3 % (4 pt
Answer #1

Let X denotes the number of residential subscribers who qualify for the favorable rates among 5 randomly selected residential subscribers.

X ~ Binomial( n = 5, p = 0.7)

The probability mass function of X is

* 0.7 * (1-0.7)5-r, x = 0, 1, 2, 3, 4, 5

a) P(X 5) = (-) * 0.75 * (1-0.7)5-5-0.75 = 0.16807

b) P(X geq 4) = P(X=4) + P(X=5)

5-4 15-5 P(X 4) 0,7(1 - 0,7

P(X > 4) 0.360 15+0.16807

P(X > 4) 0.528220

c) Here,

Mean of X = E(X) = 5*0.7 = 3.5

P( most of the residents qualify for favorable rates) = P( X > = 3.5) P(X > 4) 0.528220